OpenStudy (anonymous):

3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

i have a few others also

OpenStudy (phi):

use elimination

OpenStudy (anonymous):

can you get me started

OpenStudy (anonymous):

are you finding the value of x,y and z?? just use elimination method

OpenStudy (phi):

it might be easiest to change the order of the equations to x + 6y - z = -4 3x - 5y + z = 12 2x - 5y - 2z = -8

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation.

OpenStudy (anonymous):

the equation 1 and 2.first cancel out the z ..

OpenStudy (anonymous):

you have to use sub

OpenStudy (phi):

if we ignore the variables (but keep the same order) 1 6 -1 -4 3 -5 1 12 2 -5 -2 -8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what

OpenStudy (phi):

the first step is to make the first column below the 1 all zeros multiply the first row by -3 and add that to the 2nd row

OpenStudy (phi):

in other words, multiply -3 * 1 6 -1 -4 to get -3 -18 +3 +12 add those numbers to the 2nd row to get a new 2nd row

OpenStudy (anonymous):

these are the options a. 4x + y = 8 -17y = 0 b. 2x - 11y = 16 4x + 7y = -16 c. -23y + 4z = 24 -17y = 0 d. 13y - 2z = 24 7y - 4z = 0

OpenStudy (phi):

unless you are Rainmain, you have to go through the steps one by one

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

o-13+4+24

OpenStudy (anonymous):

or is it 6-13+4+24

OpenStudy (anonymous):

phi you there

OpenStudy (phi):

First, I am not sure where you got your numbers.

OpenStudy (anonymous):

i added both sides together

OpenStudy (phi):

2nd, looking at the "answers", I don't know what the question is ?

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x - 5y + z = 12 x + 6y - z = -4 2x - 5y - 2z = -8

OpenStudy (phi):

In that case, we are not solving for x,y,z . And they tell us to use substitution (which is a lousy way to solve for more than 2 variables... it is too confusing). But we have to use substitution.

OpenStudy (phi):

if x was isolated in the second equation. that means "solve for x" in the equation x + 6y - z = -4 can you do that ?

OpenStudy (anonymous):

yes one second

OpenStudy (anonymous):

x=-4+z-6y

OpenStudy (phi):

next, replace x with -4 + z -6y in the first equation and the 3rd equation first one: 3x - 5y + z = 12 replace x with (-4 + z -6y) 3(-4 + z -6y) - 5y + z = 12 distribute the 3 and get -12 +3z -18y -5y +z =12 can you simplify this ?

OpenStudy (anonymous):

ya one sec

OpenStudy (anonymous):

\[-12+4z-13y=12\]

OpenStudy (phi):

I would move the -12 to the other side

OpenStudy (anonymous):

oh ok

OpenStudy (phi):

and write the terms alphabetical just so it is easier to match things up

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so \[-13y+4z=24\]

OpenStudy (phi):

and double check your arithmetic on the y term

OpenStudy (anonymous):

13y not -13y

OpenStudy (phi):

?

OpenStudy (phi):

3x - 5y + z = 12 replace x with (-4 + z -6y) 3(-4 + z -6y) - 5y + z = 12 -12 +3z -18y -5y +z =12

OpenStudy (phi):

collect terms: -18y -5y +3z + z = 12+12 if you factor out y and z you get: (-18-5)y + (3+1)z = 24

OpenStudy (phi):

-18 + -5 = ?

OpenStudy (anonymous):

-23

OpenStudy (anonymous):

-23y+4z=24

OpenStudy (phi):

yes. Be careful with the arithmetic or you'll never get the correct answer, even though you understand the algebra.

OpenStudy (phi):

now see if any of the choices has that equation.

OpenStudy (anonymous):

ok i got the answer

OpenStudy (anonymous):

can you help me with another

OpenStudy (phi):

If more than one answer had that, we would do the 3rd equation. The 3rd equation turns into -17y=0 (or 17y=0 , or y=0)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

heres another

OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations. 5x + 2y + z = -2 3x + 4y + 3z = 2 -4x - 3y - 3z = 1

OpenStudy (phi):

please post it as a separate question

OpenStudy (anonymous):

ok